The number of bank accounts randomly opened per month by a Zanaco customers in Lusaka is represented by the probability distribution below.
X
19
20
21
22
P(X = x)
0.4
0.25
0.2
0.15
i. Find the expected number of accounts opened per month.
ii. Suppose opening an account incurs a fixed monthly costs of K100 and an additional handling cost of k5 per account. Find the expected monthly cost of opening bank accounts by Zanaco customers.
1
Expert's answer
2020-09-07T19:08:46-0400

solution I


Expected number of accounts per month


"\bar x = \sum x .p(x)"

"=(19*0.4)+(20*0.25)+(21*0.2)+(22*0.15)"

"=20.1"

answer: the expected number of accounts opened in a month is 20.1 or approximately 20 since you cannot create a fraction of an account


solution II


Fixed costs "=k100"

Cost per account "=k5"


Given expected monthly accounts opened "=20.1"


Expected monthly cost of opening accounts



"=k100+(20.1*k5)""=k200.5"

answer: expected monthly cost of opening accounts is k 200.5 or k 200 assuming number of accounts opened in a month are rounded off to 20 since a fraction of an account cannot be created